8 It is desired to prepare a sweetened concentrated orange juice. The initial pressed juice contains 5 % of total

solids and it is desired to lift this to 10 of total solids by evaporation and then

To add sugar to give 2% of added

Sugar in the concentrated juice.Calculate the quantity of water which must be removed, and of sugar which must

be added with respect to each 1000 kg of pressed juice

water removed 500 kg,sugar added 10.2 kg


1
Expert's answer
2021-11-08T08:27:45-0500

Initially pressed juice contains 5% of total solids which means for each 1000kg of orange juice we have 50kg (5%) of solids. Now we need to make it up to 10% 

Let's assume the water that need to be removed is Y kg then 

50kg + 2% of total solution must give 10% of total solution 

50+ 2/100 ( 100-y) = 10/100( 1000-y) 

Solving the equation gives y= 375kg 

Thus 375kg of water need to be removed to make 10% of solids 

Part 2

Sugar need to be added is 2/100(1000-375)= 12.5kg 

= 12.5kg